Joe Goodnuff had been eating on a square card table until he bought a new rectangular dining table. The new table is 2 ft longer and 1 ft wider than the card table. It has an area of 24.75 ft^2. How long was a side of his old square card table? Use the quadratic equation to solve this word problem.
Answer by amarjeeth123(569)

Let the dimensions of the square card table be x.
Then the length of the rectangular table is (x+2).
Then the width of the rectangular table is (x+1).
We have Area=Length*width
(x+2)(x+1)=24.75
x^2+3x+2=24.75
x^2+3x-22.75=0
x^2+6.5x-3.5x-22.75=0
x(x+6.5)-3.5(x+6.5)=0
(x+6.5)(x-3.5)=0
x=3.5 ft.
The side of the old square table is 3.5 feet.